# A two-circuit approach to reducing quantum resources for the quantum lattice Boltzmann method

### v2

@@ III.2 Particle streaming
The resulting statevector |$\ket{\psi_{2}}$ has a distribution shifted to a neighbor lattice site corresponding to its link value $\alpha_{i}$. The right and left shift gates $R$ and $L$, shown in fig. 4, are controlled on the link qubits $d$ to act on the part of the statevector corresponding to a distribution $\alpha$. Figure 5 shows the up and down operators $U$ and $D$, which are implemented via the $L$ and $R$ operators applied via $r_{1}$.
@@
III.3 Macroscopic variable retrieval
We retrieve the distribution $\phi(\bm{r},t)$ by summing $f_{\alpha}^{\mathrm{(eq)}}$ over the link directions $\alpha$. In terms of quantum operations, this is accomplished via the application of Hadamard gates to each of the qubits in link registers $d$, as shown in fig. 2 (c) [49]. A Hadamard gate $H$ applied to a statevector |$\ket{\psi}$, resulting in
@@ III.3 Macroscopic variable retrieval
Thus, when the Hadamard gate is applied to a link qubit in $d$, it stores the sum of the two amplitudes as the |$\ket{0}$ amplitude and the difference as the |$\ket{1}$ amplitude. So, the sum can be post-selected by ignoring |$\ket{1}$ measurements. This means that the Hadamard gates will sum the distributions but introduce a factor of $1/\sqrt{2}$ per gate. This prefactor is post-processed out of the computation by multiplication of a factor of $\sqrt{2}^{\log Q}$ where $Q=|\alpha|$ is the number of link distributions. At each time step, we retrieve the circuit result via state tomography, which must be used to extract the first $M$ lattice site elements.
IV Quantum lattice Boltzmann method for the Navier–Stokes Equations
IV.1 Stream function–vorticity formulation
@@
In this formulation, (42) and (43) use vorticity $\omega$ and stream function $\psi$ instead of directional speeds $u$ and $v$. The velocity vector is thus $\bm{u}=\{u,v\}$. The stream function relates to the directional velocities as
@@
So, (42) is a Poisson equation in stream function $\phi$ and (43) is a advection–diffusion equation in vorticity $\omega$.
IV.1.1 Lattice-based representation
@@
$\omega(\bm{r},t)$ $=\sum_{\alpha}f_{\alpha}(\bm{r},t),$ (47)
respectively. These stages, buttressed via the circuits of section III.1, enable the vorticity $\omega$ computation.
The equilibrium distribution function for the Poisson equation ($\nabla^{2}\psi=-\omega$) is $g_{\alpha}^{\text{(eq)}}(\bm{r},t)=w_{\alpha}\psi(\bm{r},t)$. The streaming and macro steps match those of section III.2 and section III.3, but the source term $S=-\omega$ is added during the collision step (a). Thus, the relaxation operator is
@@ IV.1.2 Boundary conditions
$\psi=\sum_{\alpha=0}^{N_{l}-1}g_{\alpha}^{\text{(eq)}}=g_{0}+g_{1}+g_{2}+g_{3}+g_{4},\quad\text{so,}\quad g_{\partial\Omega}=-\sum_{\alpha|\alpha\neq\partial\Omega}g_{\alpha}.$ (50)
Defining the vorticity expression (42) in terms of the stream function and expanding it in its Taylor series gives
$\omega_{i,N}=-2\left(\frac{\psi}{\Delta y^{2}}+\frac{U}{\Delta y}\right),$ (51)
@@
To implement (52), the matrix
$B=\begin{bmatrix}0&\cdots&0\\ \vdots&I_{N-2}&\vdots\\ 0&\cdots&0\end{bmatrix}$ (56)
is applied to the statevector $|\phi\rangle$, where $B$ is of size $N\times N$, where $N=\log(n)$ in each dimension and $I_{n}$ denotes the size $n$ identity matrix. So, $g_{\partial\Omega}=0$, while retaining other distribution values, enforces the boundary condition for the stream function circuit $|\psi_{\text{b.c.}}\rangle=|0\rangle^{\otimes n}$.
Due to its non-unitary properties, the vorticity circuit boundary conditions are applied via a linear combination of $B$. This linear unitary combination is
$D_{1,2}=B\pm i\sqrt{I_{N}-B^{2}},$ (57)
following section III.1.2.
Figure 6: A D2Q5 vorticity circuit, including boundary conditions. Compared with the advection–diffusion circuit in fig. 2, this has an additional qubit $b$ required to hold boundary conditions and an additional operation required to compute them. However, if boundary conditions are computed classically, the vorticity and advection–diffusion circuits are identical.
IV.2 Quantum circuits
IV.2.1 Vorticity circuit
@@
The vorticity circuit computes the vorticity $\omega$ for the current time step. Note the vorticity circuit in fig. 6 and advection–diffusion circuit in fig. 2 match, except for boundary conditions. This occurs because the vorticity equation (described in eq. 43) follows the same form as the diffusion equation. The Navier–Stokes algorithm’s vorticity circuit is identical to the advection–diffusion circuit for classically-computed boundary conditions.
@@
If boundary conditions are included, an extra qubit $b$ stores them. Boundary conditions require additional computation, as enforcing such conditions is not a unitary operation. For this, the linear combination of unitaries is used [53], described further in section III.1.2, The input to this circuit for the no-boundary version is the previous time step’s vorticity $\omega_{t-1}$. When using boundary conditions, the circuit input includes pre-computed boundary conditions in the boundary qubit $b$, computed following the descriptions in section IV.1.2.
IV.2.2 Stream function circuit
@@
The stream function circuit is the other quantum circuit used in this two-circuit model. The D2Q5 stream function circuit shown in fig. 7 is referred to for additional context. This circuit is similar in character to the advection–diffusion circuit with classical boundary conditions but an additional source. A key difference is the requirement of an extra qubit $s$ to store the source term, $S=-\omega$, as shown in fig. 7.
@@
The $C_{1}$ and $C_{2}$ gates also operate on the source term, and the Hadamard gate on qubit $s$ adds this source to the equilibrium stream function distribution function $g_{\alpha}^{\mathrm{(eq)}}$. The same qubit-addition process from section IV.2.1 applies. Boundary conditions require an additional $b$ qubit store and a boundary gate $B$, which is not unitary and again is implemented via a linear combination of matrices following (14). The previous stream function, $\psi_{t-1}$, and source term, $S=-\omega_{t-1}$, serve as the input to the stream function circuit. The boundary conditions are computed according to section IV.1.2 for our quantum boundary condition variant.
V Simulations and results
V.1 Advection–diffusion equation
@@
Example results for the QLBM circuit applied to the advection–diffusion equation are shown for two exemplar cases. In the 1D case, D1Q2 and D1Q3 lattice schemes simulate a dense concentration of $\rho=0.2$ at source $x_{i}=10$ undergoing advection and diffusion with uniform advection velocity of $c=1/5$ and diffusion coefficient $D=1/6$. The validation problem for the 2D case follows a diffusing concentration $\rho=0.3$ at source $(x_{i},y_{j})=(4,4)$ and $0.1$ elsewhere, solved via a D2Q5 scheme.
@@
Figure 8: Quantum LBM (a) D1Q2, D1Q3, and (b) D2Q5 results to the advection–diffusion equation. The initial condition in (a) is a point source of $\rho=0.2$ at $x=10$ and $\rho=0.1$ in all other sites. The initial condition in (b) is a point source of $\rho=0.3$ at $(x,y)=(4,4)$ and $\rho=0.1$ elsewhere.
Figure 8 shows the results of these simulations, showing the checkerboard pattern that arises in D1Q2 due to the distribution moving wholly into neighboring areas, resulting in half of the lattice sites having zero particles. The algorithmic deficiency is remedied via the D1Q3 lattice scheme with an $w_{0}=2/3$ of the particles, which remain in their current lattice site. Figure 8 (a) shows that the D1Q2 and D1Q3 schemes return similar results, but D1Q3 resolves the checkerboarding problem.
Figure 8 (b) shows the results of the 2D test problem solved with the D2Q5 scheme. We observe a source with an advection velocity in the positive $x$ and $y$ directions, advecting in the direction of the velocity and diffusing outwards from an initial source at $(x_{i}=4,y_{j}=4)$. The solution behaves as expected and agrees with classical LBM results.
V.2 Navier–Stokes equations
@@
Figure 9: A 2D lid-driven cavity flow. Steady-state stream function isocontours are shown as labeled. The initial conditions are $\psi=0$, $w=0$, and $U=1$.
@@
(a) Relative error $\varepsilon_{\omega}$ for vorticity circuit
(b) Relative error $\varepsilon_{\psi}$ for stream function circuit
@@
The relative errors are shown for the cavity problem in fig. 10. The two-circuit QLBM has a close agreement with the results of the classical LBM when simulated with the statevector simulator in Qiskit’s SDK [48], as shown in fig. 10. The state vector simulator perfects exact tomography but is generally limited to relatively small simulations due to exponentially increasing memory requirements with qubit numbers.
V.2.1 Quantum resource estimation and improvement
@@
Single-circuit QLBM 25 58
@@
Vorticity 12 39
Stream function without boundaries 4.2 15
Vorticity without boundaries 3 6.5
@@
Implementing the two-circuit method for solving the Navier–Stokes equations using quantum lattice-based algorithms shows quantum resource efficiency advantages over the single-circuit method. The difference in accuracy between the classical lattice Boltzmann method (LBM) and the two-circuit implementation is negligible. The stream function and vorticity circuits can be run concurrently for each time step, saving computational time. Other improvements include circuits with classical boundary conditions, decreasing circuit depth, and the number of CNOT gates compared to a single-circuit strategy. CNOT gates are slower than single qubit gates such as Hadamard gates, so CNOT gate count is used here to compare quantum resource costs.
The Qiskit transpiler translates the circuits of previous sections into single qubit and CNOT gates. The results of table 1 show that pre-computing the boundary conditions using classical methods reduces the gate count by at least $35\%$. Even when these circuits are computed in serial, the combined quantum resources for stream function and vorticity circuits remain notably lower than those of a single-circuit QLBM. Of course, parallel implementations are natural with the two-circuit approach.
VI Conclusion
This work presents improvements on the quantum lattice Boltzmann method for solving the two-dimensional Navier–Stokes equations. Testing on the validation problem shows that the error between the quantum and classical LBM methods is negligible. Runtime and quantum resources markedly decrease from a previous single-circuit implementation of an otherwise similar algorithm [42]. The number of qubits required scales as $O(\log M)$ where $M$ is the size of the grid, as opposed to the linear scaling of previous related works such as Yepez [36], which makes the prospect of near-term implementations on quantum hardware potentially viable, though not tested here.
The CNOT count of the vorticity and stream function circuits scales as $O(\alpha M)$. However, the encoding step is an efficiency bottleneck, occurring after each time step. Readout of the statevector and subsequent re-encoding decreases the efficiency of this method and is an important consideration for future work. Using the two-circuit quantum lattice Boltzmann method simplifies the quantum algorithm implementation and reduces the required quantum resources.
Acknowledgements

### v3

@@ 3.2 Particle streaming
The resulting statevector |$\ket{\psi_{2}}$ has a distribution shifted to a neighboring lattice site determined by its link value $\alpha_{i}$. Figure 4 shows how the right and left shift gates $R$ and $L$ are controlled on the link qubits $d$ to act on the part of the statevector for distribution $\alpha$. Figure 5 shows the up and down operators $U$ and $D$, which are implemented via the $L$ and $R$ operators applied via $r_{1}$.
@@
3.3 Macroscopic variable retrieval
We retrieve the distribution $\phi(\bm{r},t)$ by summing $f_{\alpha}^{\mathrm{(eq)}}$ over the link directions $\alpha$. Figure 2 shows how this is accomplished by applying Hadamard gates to each of the qubits in link registers $d$, as shown in fig. 2 (c) [80]. A Hadamard gate $H$ applied to a statevector |$\ket{\psi}$, resulting in
@@ 3.3 Macroscopic variable retrieval
When the Hadamard gate is applied to a link qubit in $d$, it stores the sum of the two amplitudes as the |$\ket{0}$ amplitude and the difference as the |$\ket{1}$ amplitude. Thus, the sum can be post-selected by ignoring |$\ket{1}$ measurements. With this post-selection, the Hadamard gates sum the distributions but introduce a factor of $1/\sqrt{2}$ per gate. This pre-factor is post-processed out of the computation by multiplication of a factor of $\sqrt{2}^{\log N_{\mathrm{links}}}$ where $N_{\mathrm{links}}=|\alpha|$ is the number of link distributions. At each time step, we retrieve the circuit result via state tomography, which must be used to extract the first $N_{\mathrm{site.}}$ lattice site elements.
4 Quantum Lattice Boltzmann method for the Navier–Stokes equations
4.1 Stream function–vorticity formulation
@@
In this formulation, 22 and 23 use vorticity $\omega$ and stream function $\psi$ instead of directional speeds $u$ and $v$. The velocity vector is thus $\bm{u}=\{u,v\}$. The stream function relates to the directional velocities as
@@
Thus, 22 is a Poisson equation in stream function $\phi$ and 23 is an advection–diffusion equation in vorticity $\omega$.
4.1.1 Lattice-based representation
@@
$\omega(\bm{r},t)$ $=\sum_{\alpha}f_{\alpha}(\bm{r},t).$ (27)
These stages, buttressed via the circuits of section 3.1, enable the vorticity $\omega$ computation.
The equilibrium distribution function for the Poisson equation ($\nabla^{2}\psi=-\omega$) is $g_{\alpha}^{\text{(eq)}}(\bm{r},t)=w_{\alpha}\psi(\bm{r},t)$. The streaming and macro steps match those of sections 3.2 and 3.3, but the source term $S=-\omega$ is added during the collision step (a). Thus, the relaxation operator is
@@ 4.1.2 Boundary conditions
$\psi=\sum_{\alpha=0}^{N_{\mathrm{links}}-1}g_{\alpha}^{\text{(eq)}}=g_{0}+g_{1}+g_{2}+g_{3}+g_{4},\quad\text{so,}\quad g_{\partial\Omega}=-\sum_{\alpha|\alpha\neq\partial\Omega}g_{\alpha}.$ (30)
Defining the vorticity expression 22 in terms of the stream function and expanding it in its Taylor series gives
$\omega_{i,N_{\mathrm{qubit}}}=-2\left(\frac{\psi}{\Delta y^{2}}+\frac{U}{\Delta y}\right),$ (31)
@@
To implement 32, the matrix
$B=\begin{bmatrix}0&\cdots&0\\ \vdots&I_{N_{\mathrm{qubit}}-2}&\vdots\\ 0&\cdots&0\end{bmatrix},$ (33)
is applied to the statevector $|\phi\rangle$, where $B$ is of size $N_{\mathrm{qubit}}\times N_{\mathrm{qubit}}$, $N_{\mathrm{qubit}}=\log(N_{\mathrm{site.}})$ in each dimension, and $I_{n}$ denotes the size $n$ identity matrix. Setting $g_{\partial\Omega}=0$, while retaining other distribution values, enforces the boundary condition for the stream function circuit $|\psi_{\text{b.c.}}\rangle=|0\rangle^{\otimes N_{\mathrm{qubit}}}$.
The vorticity circuit boundary conditions are applied via a linear combination of $B$ to account for its nonlinear nature. This linear unitary combination is
$D_{1,2}=B\pm i\sqrt{I_{N_{\mathrm{qubit}}}-B^{2}},$ (34)
following section 3.1.2.
Figure 6: The original D2Q5 vorticity circuit proposed by Budinski [50], modified from the generalized circuit in fig. 2 for boundary conditions. The changed circuit includes an additional qubit $b$ to store the boundary conditions and computations. We propose an algorithm that splits this circuit into two distinct parts.
4.2 Two-circuit model
The deviation from Budinski [52]’s original algorithm lies in the two-circuit approach, which defines distinct circuits for the stream function and vorticity computation. To devise this, we adopt the advection–diffusion circuits, as discussed in [50], and apply different bounds specific to the lid-driven cavity case. We discuss each circuit in more detail below.
4.2.1 Vorticity circuit
@@
The vorticity circuit computes the vorticity $\omega$ for the current time step. The vorticity circuit in fig. 6 and advection–diffusion circuit in fig. 2 are nearly identical, save for the boundary conditions. This matching occurs because the vorticity equation in 23 follows the same form as the diffusion equation. The Navier–Stokes algorithm’s vorticity circuit is the same as the advection–diffusion one if the boundary conditions are computed classically.
@@
If boundary conditions are included, an extra qubit $b$ stores them. Boundary conditions require additional computation, as enforcing such conditions is not a unitary operation. For this, the linear combination of unitaries is used  [85], described further in section 3.1.2. The input to this circuit for the no-boundary version is the previous time step’s vorticity $\omega_{t-1}$. When using boundary conditions, the circuit input includes pre-computed boundary conditions in the boundary qubit $b$, computed following the descriptions in section 4.1.2.
4.2.2 Stream function circuit
@@
The stream function circuit is the other quantum circuit used in this two-circuit model. The D2Q5 stream function circuit shown in fig. 7 is referred to for additional context. This circuit is similar to the advection–diffusion circuit with classical boundary conditions but includes an additional source term. Figure 7 shows the key difference between these circuits, where the required additional qubit $s$ stores the source term, $S=-\omega$.
@@
The $C_{1}$ and $C_{2}$ gates also operate on the source term, and the Hadamard gate on qubit $s$ adds this source to the stream function’s equilibrium distribution function $g_{\alpha}^{\mathrm{(eq)}}$. The same qubit-addition process from Section 4.2.1 applies. Boundary conditions require an additional $b$ qubit store and a boundary gate $B$, which is not unitary and again is implemented via a linear combination of matrices following 14. The previous stream function, $\psi_{t-1}$, and source term, $S=-\omega_{t-1}$, serve as the input to the stream function circuit. The boundary conditions are computed according to section 4.1.2 for our quantum boundary condition variant.
5 Simulations and results
The QLBM algorithm of Budinski [52] uses a single circuit to solve the Navier–Stokes equations. This work follows from corresponding work on advection–diffusion algorithms [50]. QLBM-frugal separates the computational process into distinct stream function and vorticity circuits. We verify the algorithm against that of Budinski [50]. We consider two cases for the advection–diffusion algorithm, the D1111Q2222 and D1111Q3333 lattice schemes, and validate the implementation accuracy before separating of the stream function and vorticity circuits. We also consider a lid-driven cavity problem to verify the two-circuit approach. With this, we can compare how the resources use of the algorithm scale under increasing lattice site count against the work of Budinski [50].
5.1 Advection–diffusion equation
@@
We apply the QLBM circuit to the advection–diffusion equation to obtain results for two exemplar cases. We verify these results against the expected outcomes returned by the classical LBM. In the 1D case, D1Q2 and D1Q3 lattice schemes simulate a dense concentration of $\rho=0.2$ at source $x_{i}=10$, undergoing advection and diffusion with uniform advection velocity of $c=1/5$ and diffusion coefficient $D=1/6$. The validation problem for the 2D case follows a diffusing concentration $\rho=0.3$ at source $(x_{i},y_{j})=(4,4)$ and $0.1$ elsewhere, solved via a D2Q5 scheme.
@@
Figure 8: Quantum LBM (a) D1Q2, D1Q3, and (b) D2Q5 results for the advection–diffusion equation. The initial condition in (a) is a point source of $\rho=0.2$ at $x=10$ and $\rho=0.1$ otherwise. The initial condition in (b) is a point source of $\rho=0.3$ at $(x,y)=(4,4)$ and $\rho=0.1$ elsewhere.
Figure 8 shows the results of statevector simulations over $50$ timesteps. A checkboard pattern arises in the D1Q2 case because the distribution moves wholly into neighboring areas, resulting in half of the lattice sites having zero particles. The algorithmic deficiency is remedied via the D1Q3 lattice scheme by setting the weight of vector $w_{0}$, specified in fig. 1, to $2/3$. Figure 8 (a) shows that D1Q3 resolves the checkerboarding problem of D1Q2.
Figure 8 (b) shows the results of the 2D test problem solved with the D2Q5 scheme after $20$ timesteps. There exists an initial source at $(x_{i}=4,y_{j}=4)$, with an advection velocity in the positive $x$ and $y$ directions. The source advects in the direction of the velocity and diffuses outwards. The solution behaves as expected and agrees with classical LBM results.
Figure 9: The results of the D1111Q3333 algorithm at iteration $50$ are obtained via finite sampling of the quantum circuit at various sample sizes $N_{\mathrm{samp.}}$.
The results are further validated via finite sampling. Figure 9 displays the results of finite sampling with various shot sample sizes $N_{\mathrm{samp.}}$. We verify the behavior of the algorithm by computing the state fidelity of results obtained via finite sampling to the statevector of the ideal solution. The quantum state fidelity, $F_{\psi}(\rho)$, of a (probabilistic) mixed quantum state $\rho$ with respect to a pure quantum state $\psi$ is expressed as
$F_{\psi}(\rho)=\langle\psi|\rho|\psi\rangle=\textrm{Tr}(p|\psi\rangle\langle\psi|).$ (35)
Here, $\psi$ is the expected solution obtained via the QLBM algorithm, and $\rho$ is the probabilistic outcome obtained via finite sampling methods. From the equation defined in 35, we conclude that if $\rho$ perfectly resembles the ideal solution, then we have $F_{\psi}(\rho)=1$. The fidelity is expected to be expressed as a function of the number of shots $N_{\mathrm{samp.}}$ used in the experiment, where in accordance to [86]
$N_{\mathrm{samp.}}\propto\frac{1}{1-F_{\psi}(\rho)}$ (36)
Figure 10 shows the results of this verification. The slope of the fit line in fig. 10 is $1.01$, a 1% difference from the expected proportion.
Figure 10: Verification via linear convergence with shot number of the advection–diffusion D1Q3 algorithm.
5.2 Navier–Stokes equations
@@
Figure 11: A 2D lid-driven cavity flow. Steady-state stream function isocontours are shown as labeled. The initial conditions are $\psi=0$, $w=0$, and $U=1$, done over $80$ timesteps.
@@
[Uncaptioned image] [Uncaptioned image]
(a) D1Q2 and D1Q3 (b) D2Q5
Refer to caption Refer to caption
(a) Relative error $\varepsilon_{\omega}$ for vorticity circuit (b) Relative error $\varepsilon_{\psi}$ for stream function circuit
@@
The relative errors are shown for the cavity problem in fig. 12. Figure 12 shows that the two-circuit QLBM agrees with the classical LBM when simulated with the statevector simulator in Qiskit’s SDK [79]. The statevector simulator perfects exact tomography but is generally limited to small simulations due to exponentially increasing memory requirements with qubit numbers.
5.2.1 Quantum resource estimation and improvement
@@
Single-circuit QLBM 25.0 58.0
@@
Vorticity 12.0 39.0
Stream function without boundaries 4.2 15.0
@@
Implementing the two-circuit method for solving the Navier–Stokes equations using quantum lattice-based algorithms shows quantum resource advantages over the single-circuit method. The advantages are achieved in two areas: two-qubit gate count and runtime. Two-qubit gates like CNOTs are slower and more error-prone than single-qubit gates, which has prompted bodies of work to reduce counts of these gates in quantum algorithms [87]. The projected runtime is calculated by summing the implementation time for each gate in the circuit. As such, the runtime is approximately proportional to the circuit depth.
We show the reduction in gate counts by converting the circuit into a series of equivalent one- and two-qubit gates via the Qiskit transpiler. This procedure is conducted on a $64\times 64$ lattice, with no prior optimizations or conversions to backend-specific gate sets. Table 1 show that pre-computing the boundary conditions using classical methods reduces the gate count by at least $35\%$. By running the QLBM-frugal stream function and vorticity circuits concurrently, circuit depth is reduced to the maximum depth of its constituent circuits. Table 1 shows that a 33% reduction in circuit depth when compared against the single-circuit algorithm by Budinski [52].
We similarly transpile the results relative to a select backend. During compilation, we set the optimization level to be 3, as defined in Javadi-Abhari et al. [79] documentation. The optimization levels in Qiskit range from 0 to 3, corresponding to more aggressive optimization. The most aggressive optimizations include noise adaptive qubit mapping and gate cancellation. This compilation reduces the circuit to one with fewer gates but the same functionality. Still, such optimization only modestly reduces gate counts in the cases discussed herein.
The IBM Brisbane device is the backend of the transpilation process. Brisbane supports $127$ qubits and is sufficient for resource estimation. The prior case of table 1 transpiled the circuit into a generalized set of gates. With the Brisbane backend, the transpiler identifies the device-specific gate set, which includes single-qubit X, RZ, and SX gates and the two-qubit Echoed Cross-Resonance (ECR) gate. ECR gates and a series of single-qubit rotations are applied to implement a CNOT gate. Thus, reducing the ECR gate count is commensurate with a reduced CNOT count. The transpilation, optimization, and resource estimation processes are both performed on a local classical device before simulation.
Table 2 illustrates the resource estimation for a $16\times 16$ lattice size case. All algorithms shown in the table, including the QLBM algorithm by Budinski [52], are optimized via the Qiskit transpiler before resource estimation. Table 2 shows that, after optimization, we see a 33% reduction in ECR gates. Of the total ECR gates applied, table 2 shows that the boundary conditions make up 44% of the ECR gate counts in the QLBM-frugal algorithm. It remains to be shown whether the reduction in two-qubit gates offsets the cost of computing these bounds on a classical device. Doing so involves accounting for the lattice size, hardware capabilities, and error rates of the gates. By computing the results of the stream function and vorticity circuits concurrently, the circuit depth is reduced by 41%.
ECR Gates Circuit Depth
Single-circuit QLBM 43 135
Stream function 7.6 29
Vorticity 21 80
Stream function without boundaries 7.5 28
Vorticity function without boundaries 5.0 17
Table 2: Quantum resource estimation, in units of $10^{4}$, for a D2Q5 algorithm with lattice size $16\times 16$ following transpiling with optimization level $3$ on the IBM Brisbane backend.
Figure 13: Comparison of costly two-qubit gate counts on the D2Q5 algorithm for the proposed two-circuit approach against Budinski [50]’s algorithm. Transpiled on the IBM Brisbane backend. The horizontal axis “Number of Lattice Sites” scales logarithmically.
We now consider how the resource estimation changes with respect to the lattice size. First, the growth in ECR gate counts is observed in fig. 13 as the algorithm scales from a $2\times 2$ to $32\times 32$ lattice size. Figure 13 shows the number of two-qubit gates required by Budinski [52]’s QLBM algorithm increases more rapidly than that of the proposed QLBM-frugal algorithm. The number of ECR gates is on the order of one million, implying the difference in cumulative error between the initial and proposed algorithm is notable. We obtained a runtime estimate for each transpiled circuit corresponding to these lattice sizes via the Qiskit scheduler. These are calculated based on the execution time of the gates on the specified backend. These estimates do not account for the encoding and readout costs, which are discussed in section 6.
Figure 14 shows the runtime of the respective algorithms. The stream function and vorticity circuits are presumed to run concurrently, indicating that the QLBM algorithm’s runtime scales with the fastest-growing circuit. Note that the runtime of all of the circuits considered exceeds the coherence time of the IBM Brisbane hardware, as with most available quantum hardware at present. These coherence times are on the order of several hundred microseconds. Thus, results can not be reliably computed on a real device for even the smallest number of lattice sites. Furthermore, any classical simulations that seek to mimic the error rates of available hardware will return only noise. Nonetheless, fig. 14 demonstrates that running QLBM-frugal in parallel achieves a reduction in runtime when compared to the work by Budinski [50].
Figure 14: Runtime estimation for the proposed D2Q5 QLBM-frugal algorithm without quantum bounds, compared against the QLBM algorithm proposed by Budinski [50]. The circuit is transpiled with optimization level $3$ on the IBM Brisbane backend.
6 Limitations of current work
6.1 Hardware limitations
Current quantum devices have high noise floors and low qubit counts. LBM problems are often large, requiring many qubits and so beyond the capabilities of current devices. The results here follow from quantum simulation. Concurrently executing the stream function and vorticity circuits doubles the number of required qubits if they are ”parallelized” on the same quantum device. Instead, the quantum work can be distributed across multiple quantum processors, though we limit the present work to single-device analysis.
While the qubit count is indeed doubled, the depth of each circuit is reduced from the initial algorithm proposed by Budinski [50]. This strategy reduces the overall runtime and two-qubit gate count. However, the estimated runtime exceeds the coherence time of available hardware by several orders of magnitude. The presented method still depends on quantum simulators, even for small problems.
6.2 Encoding and readout costs
The algorithm of Budinski [50] and the presented two-circuit approach begin with an arbitrary state. This approach assumes that the state has been encoded in the quantum RAM or an equivalent qubit set. Unless the amplitudes are roughly uniform, the encoding process is costly. For some quantum algorithms, this could reduce the expected quantum advantage over their classical counterparts to a polynomial one [33]. In the Qiskit implementation of state preparation, the resource cost scales exponentially with respect to the number of qubits [88]. Thus, appreciating the impacts of encoding can inform the application of QLBM algorithms.
The final readout process depends on the final solution state. Quantum state tomography obtains (reads out) the state of the quantum device. One requires $12^{d}-$ 1 measurements to determine the quantum state of a $d$-dimensional Hilbert space [89]. This exponential readout cost can meaningfully impact the runtime and required resources. This cost motivates an implementation that bypasses exponential readout costs or admits a heuristic-equipped final state for tomography. More efficient quantum state tomography methods exist in special cases, and the topic remains an active research problem [89, 90, 91, 92, 93].
7 Conclusion
This work presents improvements on the quantum lattice Boltzmann method for solving the two-dimensional Navier–Stokes equations. In particular, we propose a QLBM algorithm for concurrently circuit computation of the stream function and vorticity. These modifications give rise to the possibility of solving the Navier–Stokes equations via parallelization or distributed computing. We first verified the solution returned by the proposed QLBM-frugal algorithm against classical LBM methods. We selected the lid-driven cavity problem for algorithm verification, implementing both QLBM and QLBM-frugal using Qiskit. The error between the results is negligible on ideal quantum devices, indicating the algorithm is practical on fault-tolerant devices.
Moreover, we show QLBM-frugal achieves a reduction of the two-qubit gate count from the previous single-circuit implementation of an otherwise similar algorithm [52]. Quantum devices are bottlenecked in part by limitations due to error and environmental interference. Reducing two-qubit gates will improve the algorithm’s accuracy. We demonstrate a 33% reduction in the two-qubit ECR gate count following optimization and transpilation on the IBM Brisbane device for the $16\times 16$ lattice and a 35% reduction in CNOT gate count when the algorithm is transpiled relative to a general gate set on the larger $64\times 64$ lattice, corresponding to a reduction in the total circuit depth. Concurrent computation of the stream function and vorticity circuits reduces the depth on the $16\times 16$ lattice size case by 41%. This reduction is significant: even for small problems, nearly one million individual gates are required to implement the circuit on quantum hardware. This reduction removes $\mathcal{O}(10^{5})$ qubit rotations and gates from the computation for each iteration and reduces the accumulated error associated with continued gate application.
The reduction in ECR gates and depth extends to cases beyond the $16\times 16$ lattice size. We show that the QLBM-frugal resource requirements for lattice sizes ranging from $2\times 2$ to $32\times 32$ grow at a slower rate than that of the QLBM algorithm. The expected number of qubits scales logarithmically with the lattice size, $O(\log{M})$, slower than the linear growth demonstrated in Yepez [44]. The runtime of the present work, QLBM-frugal, scales more slowly than the traditional QLBM case. While we are far from being capable of implementing such a large circuit on current NISQ hardware, the changes make the circuit more feasible and less error prone on an ideal device relative to its predecessors.
The encoding and readout costs still remain a bottleneck for the algorithm. The processes must be performed at each iteration to input the prior state of the system and extract the information following each computation. This process is costly, and the processes scale exponentially with respect to the qubit number. The current work does not address read-in/out costs, though a full QLBM strategy on NISQ devices requires this attention.
Acknowledgments